A closely wound 0.20m x 0.20m square coil of 100 turns is exposed to an external magnetic field of 0.74T that is perpendicular to the plane of the coil. The coil is then rotated for 10s until it makes 60 ̊ with the magnetic field. Find the average induced electromotive force during the 10s interval.


1
Expert's answer
2022-04-04T17:09:30-0400

"\varepsilon=\frac \Phi t=\frac{nSB\sin\alpha}t=0.26~V."